1. Two point particles separated by 6 cm have charges q1 = 14nC and q2 = -23nC, respectively. What is the magnitude of the force that each exerts on each other?





2.Two charges q1 and q2 are separated by a distance 0.02mm in the air. The force between them is 100N. What will be the force between the q1 and q2 place 0.06mm apart in the air.


1
Expert's answer
2021-06-13T07:09:35-0400

1) "|F|= \frac{kq,q_2}{d^2}=\frac{9*10^9*14*10^{-9}*23*10^{-9}}{0.06^2}"

"=\frac{9*14*23*10^9}{0.06^2}"

"=80.5000*10^9N"

"=8.05*10^{14}N"

2) "F_1= \frac{kq_1q_2}{d^2}"

"100= \frac{9*10^9(q_1,q_2)}{(\frac{0.02}{1000})^2}"

"=|q_1q_|=444.4*10^{-20}"

"F_2=\frac{9*10^9*444.4*10^{-20}}{(\frac{0.06}{1000})^2}"

"F\n_2=11.11N"